The sum of two numbers is 40. One number is 32 more than the other. Find the two numbers?
Answer by CubeyThePenguin(3113)

x + y = 40
x = y + 32

Substitute the second equation into the first equation.

y + 32 + y = 40
2y + 32 = 40
2y = 8
y = 4

(x, y) = (36, 4)